Suppose an economy produces two products: kebabs and meat balls. In 2011, 10 kebabs are sold at $6 each, and 12 meat balls are sold at $135 each. The base year is 2011. Therefore:

Group of answer choices

A) the real GDP in 2011 is $1,680, and the GDP deflator is 100

B) the real GDP in 2011 is $1,000, and the GDP deflator is 59.5

C) the nominal GDP in 2011 is $1,680, and the GDP deflator is 59.5

D) the real GDP in 2011 is $1000, and the GDP deflator is 100
Answer #1

Answer

Real GDP is the market value of all goods and servuces produced measured at constant(Base year) Prices.

Mathematically Real GDP = sum of (Price *Quantity) of all goods wgere Prices are of base year and quantity are of current year. Here Base year is same as current year 2011. Hence we will use pboth rices and quantities of 2011.

=> Real GDP = 6*10 + 135*12 = 1680

Nominal GDP is the market value of all goods and servuces produced measured at current year Prices.

Here Current year is 2011 and base year is also 2011. Hence Real GDP = Nominal GDP

GDP deflator = (Nominal GDP / Real GDP)*100 and as Real GDP = Nominal GDP

=> GDP deflator = (Real GDP / Real GDP)*100 = 100

Hence, the correct answer is (A) the real GDP in 2011 is $1,680, and the GDP deflator is 100